# Solving High-Dimensional Partial Integral Differential Equations: The Finite Expression Method

### v1

@@ Abstract
In this paper, we introduce a new finite expression method (FEX) to solve high-dimensional partial integro-differential equations (PIDEs). This approach builds upon the original FEX and its inherent advantages with new advances: 1) A novel method of parameter grouping is proposed to reduce the number of coefficients in high-dimensional function approximation; 2) A Taylor series approximation method is implemented to significantly improve the computational efficiency and accuracy of the evaluation of the integral terms of PIDEs. The new FEX based method, denoted FEX-PG to indicate the addition of the parameter grouping (PG) step to the algorithm, provides both high accuracy and interpretable numerical solutions, with the outcome being an explicit equation that facilitates intuitive understanding of the underlying solution structures. These features are often absent in traditional methods, such as finite element methods (FEM) and finite difference methods, as well as in deep learning-based approaches. To benchmark our method against recent advances, we apply the new FEX-PG to solve benchmark PIDEs in the literature. In high-dimensional settings, FEX-PG exhibits strong and robust performance, achieving relative errors on the order of single precision machine epsilon.
@@ 2.1 Partial Integro-Differential Equation
The aim of this paper is to find the solution $u(t,x)$ for $x\in\mathbb{R}^{d}$ and $0<t<T$ to the following partial integro-differential equation [23]:
@@ 3.2 Evaluation of the Integral Term
With this estimation, evaluating $\mathbb{E}[{u}(t,x+G(x,z))]$ is much more efficient. A bound on the error from this approximation has been established for specific functions $u$ [28, 39], but further work is required to establish a bound given a general candidate ${u}$. However, numerical results indicate that the error remains within an acceptable range. While it is possible to include even higher-order terms, no improvement in performance was found, and the computational time increased significantly when computing $4^{th}$-order derivatives. Note that as $\sigma$ grows, so do the higher order terms that are not included in this truncated form. Specifically, while the odd central moments of the normal distribution are all zero, the $2n^{th}$ terms grow with $\mathcal{O}(\sigma^{2n})$. Clearly, if $\sigma\geq 1$, this may prove to be a large source of inaccuracy if the higher order derivatives of the candidate function cannot control the these terms. However, as seen in Section 4.2, numerical stability of the proposed FEX-PG (Alg. 40), up to $\sigma=1$ is observed.
@@ 3.3 Parameter Grouping
Figure 3 illustrates the application of clustering to the weights of the $i^{th}$ leaf in the tree. The clustering algorithm identifies which input dimensions should share the same coefficient, and the total number of unique coefficients can be determined by adding one to the largest value in the output vector. A new linear layer is constructed that follows the structure provided by the clustering algorithm’s output (as shown in Figure 3 (Right)). After this clustering is performed, re-learning the weights and the bias term $\beta_{i}$ from scratch ensures that no biases is carried over from the previous formulation of the solution. In practice, SciPy’s hierarchical clustering algorithm [27, 40] is used, with a threshold parameter set to $\eta:=\frac{1}{d^{2}}$. The threshold parameter determines the maximum distance cophenetic distance between elements in a cluster, in practice this hyperparameter is problem dependent.
@@ 4 Numerical Results
This section presents numerical verification of the effectiveness of the proposed methods for solving the PIDE (1). First, two 1-dimensional cases are tested in Section 4.1, followed by two high-dimensional cases (up to 100 dimensions) in Section 4.2. Since traditional methods are prone to the curse of dimensionality, the proposed FEX-PG method is compared to with the TD-based neural network (referred as TD-NN) method [23] as introduced in Section 2.1. The test cases used in ref. [23] are adopted for a true apples-to-apples comparison. In all of these examples, the same set of binary operators: “$+$”, “--” and “$\times$”, and unary operators: “00”, “$1$”, “$x$”, “$x^{2}$”, “$x^{3}$”, “$x^{4}$”, “$e^{x}$”, “$\sin(x)$” and “$\cos(x)$” are used. Along with these operators, a depth two tree structure is used, which can be seen on the far right of Figure 1.
4.1 One Dimensional PIDE
We begin the discussion by referencing two one-dimensional PIDEs. Note that in this case, there is no need to apply parameter grouping or the integral estimation discussed in Section 3. This aims to demonstrate that the vanilla FEX method is capable of solving PIDEs for one dimensional case. For the integral term, rather than applying the integral estimation from Section 3.2, a trapezoidal rule is used to calculate the integral.
@@ 4.1 One Dimensional PIDE
Here, $\nu(dz)=\lambda\phi(z)dz$, with $\phi(z)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{1}{2}\left(\frac{z-\mu}{\sigma}\right)^{2}}$ and $x\in\mathbb{R}$. The same parameters as in [23] are used. The intensity of the Poisson process is set to $\lambda=0.3$, and the jumps follow a normal distribution with $\mu=0.4$ and $\sigma=0.25$. Both $x\text{ and }t\in[0,1]$, with $T=1$. The true solution is $u(t,x)=x$.
@@
Here, $\nu(dz)=\lambda\phi(z)dz$, with $\phi(z)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{1}{2}\left(\frac{z-\mu}{\sigma}\right)^{2}}$ and $z\in\mathbb{R}$. Again, the same parameters and domain as in [23] are used: $\lambda=0.3$, $\mu=0.4$, $\sigma=0.25$, $\epsilon=0.25$ and $\theta=0$ with $x,t\in[0,1]$, and $T=1$. The true solution is $u(t,x)=x$.
@@ 4.2 High Dimensional PIDE
In this section, the newly proposed techniques in FEX, including simplifying the integral and using parameter grouping as introduced in Section 3 are adopted to extend the excellent performance of FEX from lower-dimensional cases to this high-dimensional problem.
@@
$\begin{cases}\frac{\partial u}{\partial t}+\frac{\epsilon}{2}x\cdot\nabla u(t,x)+\frac{1}{2}Tr(\sigma^{2}H(u)),\\ +\int_{\mathrm{R}^{d}}(u(t,x+z)-u(t,x)-z\cdot\nabla u(t,x))\nu(dz)=\lambda(\mu^{2}+\sigma^{2})+\theta^{2}+\frac{\epsilon}{2}||x||^{2},\\ u(T,x)=\frac{1}{d}||x||^{2}\\ \end{cases}$ (27)
@@
Figure 7b (Table 2) compares the average runtime of FEX-based methods with that of TD-NN [23] across various dimensions. While TD-NN demonstrates faster performance, FEX delivers significantly higher accuracy, making the trade-off in runtime justifiable. By introducing parameter grouping (FEX-PG,  40), the computational cost is reduced, especially in higher dimensions. This improvement is driven by two key factors: first, it becomes easier to identify a good candidate solution during the search loop, particularly in very high dimensions (80+); second, parameter grouping allows us to optimize far fewer parameters, significantly reducing the number of optimization iterations required to achieve similar accuracy. In this case, single precision accuracy within approximately 2,000 iterations of Adam is achieved, compared to as many as 20,000 iterations without parameter grouping.
@@
Due to the use of the integral estimation (27), the error introduced by the integral estimation increases with the variance $\sigma$. To assess the performance and asses the robustness of FEX-PG under higher variance scenarios, larger values of $\sigma$ are tested. Table 3 compares the relative error of the solution obtained by FEX-PG and TD-NN for PIDE (27) in $d=100$ dimensions across various levels of variance. Even with a variance of $1$, FEX-PG performs exceptionally well, maintaining a relative error on the order of $10^{-7}$, whereas the error of TD-NN increases as the variance grows. It is important to note that, since the integral term is approximated using a Taylor series with respect to the random variable $z$, the use heavy-tailed distributions for $z$ may cause instability, due to the the need of finite moments (without this, the remainder of the Taylor series may not converge). While this limitation may inspire future research, it remains unexplored for the time being.
@@ 4.3 Conclusion and Discussion
This paper presents a numerical approach to solving the PIDE using a new finite expression method, the FEX-PG. While the standard finite expression method works well for solving one-dimensional PIDEs, its efficiency decreases significantly during the search loop and parameter optimization as the dimensionality increases. To address this, two improvements are proposed: first, simplifying the estimation of the integral term, and second, grouping trainable parameters in the candidate solution to reduce the total number of parameters that need to be optimized. Through extensive experiments on problems with various dimensions (up to 500), the method’s effectiveness in accurately identifying the true solution and achieving machine-level precision is demonstrated. In contrast, the baseline neural network approach struggles with higher dimensions and variance, a key parameter in the PIDE.
@@ Acknowledgement
Y. H. was partially supported by the US National Science Foundation under awards DMS-2244988, DMS-2206333, the Office of Naval Research Award N00014-23-1-2007, and the DARPA D24AP00325-00.

### v2

@@ Abstract
Partial integro-differential equations (PIDEs) have broad applications in the sciences, from electro-magnetism to options pricing. In this paper, we introduce a new finite expression method (FEX) to solve PIDEs. This approach builds upon the original FEX and its inherent advantages with new advances: 1) A novel method of parameter grouping is proposed to reduce the number of coefficients in high-dimensional function approximation; 2) A Taylor series approximation method is implemented to significantly improve the computational efficiency and accuracy of the evaluation of the integral terms of PIDEs. The new FEX based method, denoted FEX-PG to indicate the addition of the parameter grouping (PG) step to the algorithm, provides both high accuracy and interpretable numerical solutions, with the outcome being an explicit equation that facilitates intuitive understanding of the underlying solution structures. These features are often absent in traditional methods, such as finite element methods (FEM) and finite difference methods, as well as in deep learning-based approaches. To benchmark our method against recent advances, we apply the new FEX-PG to solve benchmark PIDEs in the literature. In high-dimensional settings, FEX-PG exhibits strong and robust performance, achieving relative errors on the order of single precision machine epsilon, significantly outperforming existing approaches based on neural networks.
@@ 2.1 Partial Integro-Differential Equation
The primary aim of this paper is to find the solution $u(t,x)$ for $x\in\mathbb{R}^{d}$ and $0<t<T$ to the following partial integro-differential equation [25]:
@@ 3.2 Evaluation of the Integral Term
Note that the term involving the gradient disappears when we take the expectation as $\mathbb{E}[z-\mu]=\mathbb{E}[z-\mathbb{E}[z]]=0$ by linearity of the expectation. With this estimation, evaluating $\mathbb{E}[{u}(t,x+G(x,z))]$ is much more efficient. A bound on the error from this approximation has been established for specific functions $u$ [30, 41], but further work is required to establish a bound given a general candidate ${u}$. However, numerical results indicate that the error remains within an acceptable range. While it is possible to include even higher-order terms, no improvement in performance was found, and the computational time increased significantly when computing $4^{th}$-order derivatives. Note that as $\sigma$ grows, so do the higher order terms that are not included in this truncated form. Specifically, while the odd central moments of the normal distribution are all zero, the $2n^{th}$ terms grow with $\mathcal{O}(\sigma^{2n})$. Clearly, if $\sigma\geq 1$, this may prove to be a large source of inaccuracy if the higher order derivatives of the candidate function cannot control the these terms. However, as seen in Section 4.3, numerical stability of the proposed FEX-PG (Alg. 40), up to $\sigma=1$ is observed.
@@ 3.3 Parameter Grouping
Figure 3 illustrates the application of clustering to the weights of the $i^{th}$ leaf in the tree. The clustering algorithm identifies which input dimensions should share the same coefficient, and the total number of unique coefficients can be determined by adding one to the largest value in the output vector. A new linear layer is constructed that follows the structure provided by the clustering algorithm’s output (as shown in Figure 3 (Right)). After this clustering is performed, re-learning the weights and the bias term $\beta_{i}$ from scratch ensures that no biases is carried over from the previous formulation of the solution. In practice, SciPy’s hierarchical clustering algorithm [29, 42] is used, with a threshold parameter set to $\eta:=\frac{1}{d^{2}}$. The threshold parameter determines the maximum cophenetic distance between elements in a cluster, in practice this hyperparameter is problem dependent.
@@ 4 Numerical Results
This section presents numerical verification of the effectiveness of the proposed methods for solving the PIDE (1). First, two 1-dimensional cases are tested in Section 4.1, followed by two high-dimensional cases (up to 100 dimensions) in Section 4.3. Since traditional methods are prone to the curse of dimensionality, the proposed FEX-PG method is compared to with the TD-based neural network (referred as TD-NN) method [25] as introduced in Section 2.1. The test cases used in ref. [25] are adopted for a true apples-to-apples comparison. In all of these examples, the same set of binary operators: “$+$”, “--” and “$\times$”, and unary operators: “00”, “$1$”, “$x$”, “$x^{2}$”, “$x^{3}$”, “$x^{4}$”, “$e^{x}$”, “$\sin(x)$” and “$\cos(x)$” are used. Along with these operators, a depth-2 tree structure is used, which can be seen on the far right of Figure 1.
4.1 Low Dimensional PIDE
We begin the discussion by referencing two one-dimensional PIDEs from [25]. Note that in this case, there is no need to apply parameter grouping or the integral estimation discussed in Section 3. This aims to demonstrate that the vanilla FEX method is capable of solving PIDEs for one dimensional case. For the integral term, rather than applying the integral estimation from Section 3.2, a trapezoidal rule is used to calculate the integral. No stopping criteria other than a maximum number of iterations was used for these examples as they serve to demonstrate vanilla FEX’s potential at solving PIDEs.
@@ 4.1 Low Dimensional PIDE
Here, $\nu(dz)=\lambda\phi(z)dz$, with $\phi(z)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{1}{2}\left(\frac{z-\mu}{\sigma}\right)^{2}}$ and $x\in\mathbb{R}$. The same parameters as in [25] are used. The intensity of the Poisson process is set to $\lambda=0.3$, and the jumps follow a normal distribution with $\mu=0.4$ and $\sigma=0.25$. Both $x\text{ and }t\in[0,1]$, with $T=1$. In practice, it was found that using 50 evenly spaced points in $[0,1]$ for numerical integration (via the trapezoid rule) was sufficient. The true solution of Eqn. (25) is $u(t,x)=x$.
@@
Here, $\nu(dz)=\lambda\phi(z)dz$, with $\phi(z)=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{1}{2}\left(\frac{z-\mu}{\sigma}\right)^{2}}$ and $z\in\mathbb{R}$. Again, the same parameters and domain as in [25] are used: $\lambda=0.3$, $\mu=0.4$, $\sigma=0.25$, $\epsilon=0.25$ and $\theta=0$ with $x,t\in[0,1]$, and $T=1$. Once again, 50 points evenly spaced in $[0,1]$ were sufficient for numerical integration. The true solution is $u(t,x)=x$.
@@
4.2 The Variance Gamma Model
To further demonstrate our method’s strength, we present another example. While in low dimension, this problem shows FEX’s strength at solving problems beyond toy examples and function approximation. In addition, we use this problem to verify the proposed integral simplification method given in Section 3.2. Motivated by [10], we seek to solve the Variance Gamma (VG) options pricing problem. The VG model is a widely used pure-jump Lévy process that captures the observed features of asset returns that are not well modeled by the classical Black-Scholes framework. In the context of European option pricing, the VG model leads to a PIDE. Letting $V(S,t)$ be the value of the option given asset price $S$ and time until option maturity $t$ the PIDE governing the model’s behavior can be written as
[…]
The measure $k(y)$ is the Lévy density of the VG process, which is given by
$k(y)=\frac{e^{-\lambda_{p}y}}{\nu y}1_{y>0}+\frac{e^{-\lambda_{n}|y|}}{\nu|y|}1_{y<0},$ (28)
where
$\lambda_{p}=\Big{(}\frac{\theta^{2}}{\sigma^{4}}+\frac{2}{\sigma^{2}\nu}\Big{)}^{\frac{1}{2}}-\frac{\theta}{\sigma^{2}},$ (29)
and
$\lambda_{n}=\Big{(}\frac{\theta^{2}}{\sigma^{4}}+\frac{2}{\sigma^{2}\nu}\Big{)}^{\frac{1}{2}}+\frac{\theta}{\sigma^{2}}.$ (30)
We choose $\theta=-0.4$, $\sigma=0.4$, $\nu=0.4$, $r=0.05$ and $q=0.02$, all as in [10]. To facilitate training we rescale the problem so that the asset value $S$ lies within $[0,1]$. Our strike price is likewise scaled, and is then $K=1/3$. Time to maturity, $t$ is within $[0,2]$. Note that $r$ is the risk-free interest rate and $q$ is the dividend rate of the stock. We will model European puts, so our initial condition is $V(0,S)=(S-K)^{+}$ where $(S-K)^{+}=\max(S-K,0)$. Because of this initial condition, another change made is to include ReLU as a unary function that FEX can use to represent the solution - this choice is straightforward given the form of $\max(S-K,0)$, and will allow a better fit while keeping expression size short. The boundary conditions are given as $V(t,0)=Ke^{-rt}$ and $V(t,1)=0$. Note that we are not making the common change of variables $x=log(S)$. Rather than use log-price as our variable we simply use raw asset price. The functional $\mathcal{L}$ (17) now includes an extra term as we have both boundary conditions and an initial condition. Letting $\mathcal{D}$ once again represent the PIDE, in this case Eqn. (27), we have:
$\mathcal{L}(u)\approx\frac{1}{N}\sum_{i=1}^{N}|\mathcal{D}(\tilde{u}(t_{i},S_{i}))|^{2}+\frac{1}{M}\sum_{j=1}^{M}|\tilde{u}(t_{j},S_{min})-Ke^{-rt}|^{2}+|\tilde{u}(t_{j},S_{max})|^{2}+|\tilde{u}(0,S_{j})-(S_{j}-K)^{+}|^{2}.$ (31)
As noted earlier, to evaluate the integral term we employ the same Taylor approximation trick as proposed in Section 3.2, with the slight difference being that we need to find $\mathbb{E}[{u}(t,xG(x,y))]=\mathbb{E}[{u}(t,xe^{y})]$ rather than $\mathbb{E}[{u}(t,x+z)]$.
To benchmark our results, we will compare FEX’s results to those from the Carr-Madan Fast Fourier transform based method [3]. It is important to note that an explicit solution to this problem does not exist. Hence FEX is learning an approximation, which greatly increases the challenge.
Figure 7: Plot of the put value surfaces given by the Carr-Madan FFT method [3], FEX, and the difference between the two.
The true solution, the numerical solution, and their difference is visulaized in Figure 7. The average mean square error (measured across ten consecutive runs of FEX) between FEX’s solution and the Carr-Madan solution was $2.25\times 10^{-4}$, indicating a good fit. 300 iterations of the searching loop were performed using a depth-3 tree (one level deeper than the tree seen on the far right of Figure 1). Due to the lack of an explicit true solution we did not use a stopping criteria based on accuracy; rather, we simply used a fixed number of fine-tune steps (30,000). Importantly, no significant change in accuracy was found when numerical integration (using the trapezoid rule) was performed instead of the simplification discussed in Section 3.2, but the speedup was very significant.
@@ 4.3 High Dimensional PIDE
Having established the strength of FEX in solving PIDEs and functional approximation in lower dimensions, all newly proposed techniques in FEX, including simplifying the integral and using parameter grouping as introduced in Section 3,are adopted to extend the excellent performance of FEX from lower-dimensional cases to this high-dimensional problem.
@@
$\begin{cases}\frac{\partial u}{\partial t}+\frac{\epsilon}{2}x\cdot\nabla u(t,x)+\frac{1}{2}Tr(\sigma^{2}H(u)),\\ +\int_{\mathrm{R}^{d}}(u(t,x+z)-u(t,x)-z\cdot\nabla u(t,x))\nu(dz)=\lambda(\mu^{2}+\sigma^{2})+\theta^{2}+\frac{\epsilon}{2}||x||^{2},\\ u(T,x)=\frac{1}{d}||x||^{2}.\\ \end{cases}$ (32)
@@
Figure 8b (Table 2) compares the average runtime of FEX-based methods with that of TD-NN [25] across various dimensions. For all experiments, including these timed high dimensional runs, FEX-PG was run on a single NVIDIA A100 80G. As seen in Algorithm 40, threshold parameter $\eta=\frac{1}{d}$ where $d$ is the number of space dimensions( i.e. $x\in\mathrm{R}^{d}$). 50 iterations of the searching loop were performed, and the fine-tuning was stopped when the loss functional $\mathcal{L}$ was below $1.5\times 10^{-14}$. The cut off (of near to double precision epsilon) corresponded well with the found solution having a relative error near to single precision epsilon. This allowed us to investigate how much time was saved with the parameter grouping step while holding the accuracy of the found solution largely constant, as seen in Figure 8a and Figure 8b respectively. While TD-NN demonstrates faster performance, FEX delivers significantly higher accuracy, making the trade-off in runtime justifiable. By introducing parameter grouping (FEX-PG,  40), the computational cost is reduced, especially in higher dimensions. This improvement is driven by two key factors: first, it becomes easier to identify a good candidate solution during the search loop, particularly in very high dimensions (80+); second, parameter grouping allows us to optimize far fewer parameters, significantly reducing the number of optimization iterations required to achieve similar accuracy. In this case, single precision accuracy within approximately 2,000 iterations of Adam is achieved, compared to as many as 20,000 iterations without parameter grouping.
@@
Due to the use of the integral estimation (32), the error introduced by the integral estimation increases with the variance $\sigma$. To assess the performance and asses the robustness of FEX-PG under higher variance scenarios, larger values of $\sigma$ are tested. Table 3 compares the relative error of the solution obtained by FEX-PG and TD-NN for PIDE (32) in $d=100$ dimensions across various levels of variance. Even with a variance of $1$, FEX-PG performs exceptionally well, maintaining a relative error on the order of $10^{-7}$, whereas the error of TD-NN increases as the variance grows. This good behavior of FEX is also demonstrated by the earlier example of the VG model, where the variance of $k(y)$, the Lévy density of the VG process is $0.064$. It is important to note that, since the integral term is approximated using a Taylor series with respect to the random variable $z$, the use of heavy-tailed distributions for $z$ may cause instability, due to the the need of finite moments (without this, the remainder of the Taylor series may not converge). While this limitation may inspire future research, it remains unexplored for the time being.
@@ 4.4 Conclusion and Discussion
This paper presents a numerical approach to solving the PIDE using a new finite expression method, the FEX-PG. While the standard finite expression method works well for solving low-dimensional PIDEs, its efficiency decreases significantly during the search loop and parameter optimization as the dimensionality increases. To address this, two improvements are proposed: first, simplifying the estimation of the integral term, and second, grouping trainable parameters in the candidate solution to reduce the total number of parameters that need to be optimized. Through extensive experiments on problems with various dimensions (up to 500), the method’s effectiveness in accurately identifying the true solution and achieving machine-level precision is demonstrated. In contrast, the baseline neural network approach struggles with higher dimensions and variance, a key parameter in the PIDE.
@@ Acknowledgement
S. W. was partially supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research SciDAC program under Contract No. DE-AC02-05CH11231. Y. H. was partially supported by the US National Science Foundation under awards DMS-2244988, DMS-2206333, the Office of Naval Research Award N00014-23-1-2007, and the DARPA D24AP00325-00. Approved for public release; distribution is unlimited.